I need help solving the following problem.

Solve the system using elimination.

2x+5y+3z = -20

3x+ 2y -4z = 4

3x -3y +2z = -20

x=_______________

y=________________

z=_________________
Answer #1

criven system, 2.2 + 5y +32 = -20 3x+2y-42 = 4 3x - 3y +2Z = -20 Rewrite the system in madrix by Craussian Elimination : 1 2*18: 29 ta roles T- 01-417Ri= Rit 26237 1 olot R2 = R2-17 R3 ) 0 1 -4 0 =-4 ₹ 76 y zo